# Conditional Wasserstein Distances with Applications in Bayesian OT Flow Matching

### v1

@@ 5 Geodesics and velocity fields
    The curve $(\mu_{t})_{y}:=(1-t)\pi^{1}+t\pi^{2})_{\sharp}\alpha_{y,y}$ is the disintegration of $\mu_{t}$ with respect to $P_{Y}$. Further, $(\mu_{t})_{y}$ is a geodesic in $(\mathcal{P}_{2}(\mathbb{R}^{m}),W_{2})$ for $P_{Y}$-a.e. $y\in\mathbb{R}^{d}$.
@@
By the following proposition, the above geodesic $\mu_{t}$ has an associated vector field $v_{t}$ which satisfy a continuity equation. Moreover, informally speaking, the associated vector field $v_{t}$ does not transport any mass in the $y$-component. This is related to the observation in [18].
@@
Let $\mu_{0},\mu_{1}\in\mathcal{P}_{2,Y}(\mathbb{R}^{d}\times\mathbb{R}^{m})$. Let $\alpha\in\Gamma_{Y}^{4}(\mu_{0},\mu_{1})$ be an optimal plan in (9) and $\mu_{t}=(e_{t})_{\sharp}\alpha$, $t\in[0,1]$. Then there exists a vector field $v_{t}\in L^{2}_{\mu_{t}}(\mathbb{R}^{d}\times\mathbb{R}^{m},\mathbb{R}^{d}\times\mathbb{R}^{m})$ such that the following relations are fulfilled:
@@
    $(e_{t})_{\sharp}\left((y_{2},x_{2})-(y_{1},x_{1})\right)\alpha)=v_{t}\mu_{t}$,
@@
    for $j\leq d$, we have $(v_{t})_{j}=0$ $\mu_{t}$-a.e.,
@@
    in a distribultional sense, i.e. we have for all $\varphi\in C^{\infty}_{c}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ that
    $\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\varphi\,\mathrm{d}\mu_{t}=\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\langle\nabla\phi,v_{t}\rangle\,\mathrm{d}\mu_{t}.$ (17)
@@
Furthermore, since by Lemma 5 iii), a geodesic induced by an optimal $W_{2,Y}$ plan is weakly continuous, we obtain the following proposition from [6, Proposition 8.1.8] which is needed in the numerical section.
@@
For special cases we can drop the requirements (18) on the Borel vector field as the following proposition, which is proved in the appendix, shows.
@@
A Benamou-Brenier type formula for $W_{p,Y}$ is given in Appendix F.
@@ 6 Relaxation of the conditional Wasserstein distance
    We cannot use standard optimal transport algorithms [20] out of the box.
@@
    Assume that we are interested in the optimal plan $\alpha\in\Gamma_{Y}^{4}(\mu_{0},\mu_{1})$, but we are only given empirical measures $\mu_{n,0},\mu_{n,1}$, which are $W_{2}$ approximations of $\mu_{0},\mu_{1}$. Let $Y_{n}$ be a random variable distributed as $\pi^{1}_{\sharp}\mu_{n,0}$. Even if we assume that $\pi^{1}_{\sharp}\mu_{n,0}=\pi^{1}_{\sharp}\mu_{n,1}$, Example 9 shows that we cannot guarantee that there exists a sequence of the optimal plans $\alpha_{n}\in\Gamma_{Y_{n}}^{4}(\mu_{n,0},\mu_{n,1})$ that converges in any sense to $\alpha$.
@@
We consider independent, standard normal distributed random variables $Y,X,Z\in\mathbb{R}$. Let $\mu=\nu:=P_{Y,X}$. Now we sample $(y_{i},x_{i},z_{i})\sim(Y,X,Z)$ and define
@@
i.e. $\mu_{n}\to\mu$ and $\nu_{n}\to\nu$ as $n\to\infty$ in the $W_{2}$-topology. Let $Y_{n}$ be a random variable distributed like $\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i}}$. Then $\Gamma_{2,Y_{n}}(\mu_{n},\nu_{n})$ contains exactly one plan $\alpha_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i},x_{i},y_{i},z_{i}}$ which is clearly optimal. Let $\Delta:\mathbb{R}^{3}\to\mathbb{R}^{4}$ be defined by $\Delta(y,x,z)=(y,x,y,z)$. Then $\hat{\alpha}:=\lim_{n}\alpha_{n}=\Delta_{\sharp}\left(P_{Y}\otimes P_{X}\otimes P_{Z}\right)$ in the $W_{2}$-topology. Moreover, $\hat{\alpha}\in\Gamma_{2,Y}(\mu,\nu)$ and
@@
In order to overcome the above drawbacks, we define a cost function for which the OT plan $\alpha\in\Gamma(\mu_{o},\mu_{1})$ only approximately fulfills $\pi^{1,3}_{\sharp}\alpha=\Delta_{\sharp}P_{Y}$:
@@
The same idea has been pursued in [2], where the authors rescaled the $y$-costs to obtain a blocktriangular map in the Knothe-Rosenblatt setting [33, 46] and similarly in [29]. Note that [2] was published on ArXiv after our first version of the present paper.
@@
The following proposition shows that the issue raised in Example 9 is addressed by $W_{2,d_{\beta}}$, at least on compact sets.
@@
Assume that $\mu,\nu\in P_{2,Y}(A\times B)$ for compact subsets $A\subset\mathbb{R}^{d},B\subset\mathbb{R}^{m}$ and let $\mu_{n},\nu_{n}$ be empirical measures that converge in $W_{2}$ to $\mu,\nu$. Then for a sequence $\beta_{k}\to\infty$ there exists an increasing subsequence $n_{k}$ and optimal plans $\alpha_{n_{k}}\in\Gamma(\mu_{n_{k}},\nu_{n_{k}})$ for $W_{2,d_{\beta_{k}}}(\mu_{n_{k}},\nu_{n_{k}})$ such that $\alpha_{n_{k}}$ converges with respect to $W_{2}$ to an optimal plan $\alpha\in\Gamma^{4}_{Y}(\mu,\nu)$ for $W_{2,Y}(\mu,\nu)$.
@@ 7 Bayesian flow matching
In this section, we combine the conditional Wasserstein distance with Bayesian flow matching. We start by briefly recalling flow matching and its OT variant. Then we turn to the conditional setting, where we describe a method from the literature, which we call "diagonal" Bayesian flow matching and introduce our new OT Bayesian flow matching.
@@
The objective $L_{FM}$ can be also derived differently, with ideas inspired by the score matching framework [30, 52]. Then instead of regressing to the true velocity field at $x_{t}$, they show that regressing to it has the same gradients when one conditions at $x$ ("conditional" flow matching [35, Theorem 2]), which leads to the same loss formulation.
@@
Diagonal Bayesian Flow Matching
@@
This yields the diagonal Bayesian flow matching loss
$L_{Y,\text{FM}}(\theta)=\mathbb{E}_{(z,y,x)\sim P_{Z}\otimes P_{Y,X},t\sim U([0,1])}[\|v_{t}^{\theta}(y,x_{t})-(x-z)\|^{2}].$
Under the assumption $y_{i}\neq y_{j}$ for $i\neq j$ this diagonal matching coincides with the only admissible plan in the conditional Wasserstein distance. In general however, according to Example 9, this approach does not approximate OT plans as $X$ and $Z$ are essentially independently.
@@
Given a minibatch $(\bm{z},\bm{y},\bm{x})=\left((z_{i},y_{i},x_{i})\right)_{i=1}^{I}$ sampled from the product distribution $P_{Z}\otimes P_{Y,X}$, we can calculate the optimal map $T$ and corresponding conditional coupling $\alpha$ between $(\bm{y},\bm{z})$ and $(\bm{y},\bm{x})$. The plan $\alpha$ by construction is only supported on $(y_{i},z_{i},y_{i},(\pi^{2}\circ T)(y_{i},x_{i})))$. Now let $(y,z,y,x)\sim\alpha$, then we have
@@
$L_{Y,\text{OT}}(\theta)=\mathbb{E}_{((y,z,y,x)\sim\alpha,t\sim U([0,1])}[\|v^{\theta}_{t}(y,x_{t})-(x-z)\|^{2}].$
@@ 8.1 Particle flow convergence
In this example, we minimize $W_{Y,\beta}(P_{Y,X},P_{Y,Z})$ for the empirical measures We consider the particle flow, i.e., the flow from $(Y,Z)$ to $(Y,X)$ for empirical distributions, where we compute a particle flow [4] from $(Y,Z)$ to $(Y,X)$. We follow a particle flow path, i.e., we define a curve $z(0)_{i}\sim\mathcal{N}(0,I)$ which follows
@@
for an appropriate "distance" $\mathcal{D}$, step size $\eta$ and given joint samples $(y_{i},x_{i})_{i=1}^{n}$. We choose $\mathcal{D}$ as an approximation of $W_{2,\beta}$ by rescaling $Y$ and using the Sinkorn divergence as the sample based distance measure [21, 19], where the "blur" parameter is chosen so small that it is close to the Wasserstein distance. This way we can numerically verify the convergence in Proposition 10. Note that there are no neural networks involved in this example.
We see in Fig. 1 that increasing $\beta$ indeed yields plans which transport no mass in $Y$-direction anymore, which has the consequence that the generated images fit the corresponding label. It can be seen that for $\beta=5$ each row only has one type of digit.
@@ 8.2 GMM example
Here we use an experimental setup from [26]: in (1), we choose $P_{X}$ to be a GMM in $\mathbb{R}^{5}$ with 10 mixture components, uniformly chosen means in $[-1,1]$ and standard deviation 0.1. We apply linear diagonal forward model $f=(f_{i,j})_{i,j=1}^{5}$ with $f_{i,i}=\frac{0.1}{i+1}$ and zero components otherwise and Gaussian noise with standard deviation $0.1$. This yields a posterior measure $P_{X|Y=y}$ which is a also a GMM [26, Lemma 11]. Therefore we can sample and evaluate the true posterior as groundtruth. We train a diagonal Bayesian flow matching model according to $L_{Y,FM}$ and our OT Bayesian flow matching according to $L_{Y,OT}$ with POT [20] for the same amount of time on a fixed dataset of size $10000$, where we choose the best model according to a validation set of size $2000$. We parameterize both models with around 140k parameters. We evaluate them using the Sinkhorn distance [21, 19] with the package GeomLoss averaged with 100 posteriors and over 10 training and test runs with randomly sampled mixtures. The sampling is done via an explicit Euler discretization of 10 steps. Our proposed model trained according to $L_{Y,OT}$ with $\beta=20$ obtains an average Sinkhorn distance of 0.0235, whereas the naive model obtains a value of 0.0255. In Figure 2 one can see that both models approximate the posterior very well.
Figure 2: Posterior histograms for different methods with diagonal Bayesian flow matching on the left and our OT Bayesian flow matching on the right. Ground truth posterior is in orange and model prediction in blue.
@@ 8.3 Class conditional image generation
FID (AD) 5.3 5.1 4.79 4.57 4.38 5.34
FID (Euler) 11.18 11.05 10.82 10.13 9.98 10.76
Figure 3: Class Conditional CIFAR results for different choices of $\beta$ and for the diagonal couplings. Additionally FID results are reported using an adaptive step size solver (AD) and an Euler scheme with 20 steps (Euler).
We apply our Bayesian OT flow matching for conditional image generation. We choose the condition $Y$ to be the class labels in order to generate samples of CIFAR10 for a given class. We simulate the flows for different values of $\beta$, by which we mean that we rescale the $Y$-component by $\beta$ as mentioned in Remark 11. We also simulate flows using the "diagonal" plans which coincide with the diagonal Bayesian flow matching objective [53]. For high quality inference we use an adaptive step size solver (Runge-Kutta of order 5). We also compare quantitatively when sampling with a Euler scheme with 20 steps. The samples in Fig. 3 are generated using the adaptive step size solver and sorted by class labels. For low values of $\beta$ we see that the resulting samples do not match their class labels, increasing $\beta$ leads to accurate class representations. The samples are generated given the labels of the training samples, therefore we see improved FID results as $\beta$ increases. The diagonal flow matching objective has the correct class representations, however since the associated couplings are not optimal our experiments suggest that this leads to higher variance during training and therefore slightly lower image quality, see [50] for more details on the advantages of OT based flow matching. We evaluate the methods on a fixed number of epochs, for completeness we note that the diagonal method improves to an FID (AD) of 4.88 under additional training.

### v2

@@ 5 Geodesics and velocity fields
    The curve $(\mu_{t})_{y}:=(1-t)\pi^{1}+t\pi^{2})_{\sharp}\alpha_{y,y}$ is a disintegration of $\mu_{t}$ with respect to $P_{Y}$. Further, $(\mu_{t})_{y}$ is a geodesic in $(\mathcal{P}_{2}(\mathbb{R}^{m}),W_{2})$ for $P_{Y}$-a.e. $y\in\mathbb{R}^{d}$.
@@
By the following proposition, the above geodesic $\mu_{t}$ has an associated vector field $v_{t}$ such that $(\mu_{t},v_{t})$ satisfy a continuity equation. Moreover, informally speaking, the associated vector field $v_{t}$ does not transport any mass in the $y$-component. This is related to the observation in [20, Section 4.2].
@@
Let $\mu_{0},\mu_{1}\in\mathcal{P}_{2,Y}(\mathbb{R}^{d}\times\mathbb{R}^{m})$. Let $\alpha\in\Gamma_{Y}^{4}(\mu_{0},\mu_{1})$ be an optimal plan in (9) and $\mu_{t}=(e_{t})_{\sharp}\alpha$, $t\in[0,1]$. Then there exists a Borel measurable vector field $v:[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}\to\mathbb{R}^{d}\times\mathbb{R}^{m},\,v(t,y,x)=v_{t}(y,x)$ with $v_{t}\in L^{2}_{\mu_{t}}(\mathbb{R}^{d}\times\mathbb{R}^{m},\mathbb{R}^{d}\times\mathbb{R}^{m})$ for a.e. $t\in[0,1]$ such that the following relations are fulfilled:
@@
    $(e_{t})_{\sharp}\left((y_{2},x_{2})-(y_{1},x_{1})\right)\alpha)=v_{t}\mu_{t}$ for a.e. $t\in[0,1]$,
@@
    for $j\leq d$ we have that $v_{j}=0$ for all $(t,y,x)\in[0,1]\times\mathbb{R}^{d}\times\mathbb{R}^{m}$,
@@
    in a distributional sense, i.e. we have for all $\varphi\in C^{\infty}_{c}((0,1)\times\mathbb{R}^{d}\times\mathbb{R}^{m})$ that
    $\int_{0}^{1}\int_{\mathbb{R}^{d}\times\mathbb{R}^{m}}\frac{\partial}{\partial t}\varphi+\langle\nabla_{y,x}\phi,v_{t}\rangle\,\mathrm{d}\mu_{t}\mathrm{d}t=0.$ (17)
Here $C_{c}^{\infty}$ denotes the smooth functions with compact support.
@@
Furthermore, since by Lemma 5 iii), a geodesic induced by an optimal $W_{2,Y}$ plan is weakly continuous, we obtain the following proposition from [6, Proposition 8.1.8] which gives a connection to a flow equation and is needed for flow matching.
@@
For special cases we can drop the requirement (18) on the Borel vector field as the following proposition, which is proved in the Appendix D, shows.
@@ 6 Relaxation of the conditional Wasserstein distance
    We cannot use standard optimal transport algorithms like [22] out of the box.
@@
    Assume that we are interested in the optimal plan $\alpha\in\Gamma_{Y}^{4}(\mu_{0},\mu_{1})$, but we are only given empirical measures $\mu_{n,0},\mu_{n,1}$, which are $W_{2}$ approximations of $\mu_{0},\mu_{1}$. Let $Y_{n}$ be a random variable distributed as $\pi^{1}_{\sharp}\mu_{n,0}$. Even if we assume that $\pi^{1}_{\sharp}\mu_{n,0}=\pi^{1}_{\sharp}\mu_{n,1}$, Example 9 shows that we cannot guarantee that there exists a sequence of the optimal plans $\alpha_{n}\in\Gamma_{Y_{n}}^{4}(\mu_{n,0},\mu_{n,1})$ that converges in any meaningful sense to $\alpha$.
@@
We consider independent, standard normally distributed random variables $Y,X,Z\in\mathbb{R}$. Let $\mu=\nu:=P_{Y,X}$. Now we sample $(y_{i},x_{i},z_{i})\sim(Y,X,Z)$ and define
@@
i.e. $\mu_{n}\to\mu$ and $\nu_{n}\to\nu$ as $n\to\infty$ in the $W_{2}$-topology. Let $Y_{n}$ be a random variable distributed like $\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i}}$. Then with probability one $\Gamma_{Y_{n}}^{4}(\mu_{n},\nu_{n})$ contains exactly one plan $\alpha_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i},x_{i},y_{i},z_{i}}$ which is clearly optimal. Let $\Delta:\mathbb{R}^{3}\to\mathbb{R}^{4}$ be defined by $\Delta(y,x,z)=(y,x,y,z)$. Then $\hat{\alpha}:=\lim_{n}\alpha_{n}=\Delta_{\sharp}\left(P_{Y}\otimes P_{X}\otimes P_{Z}\right)$ in the $W_{2}$-topology. Moreover, $\hat{\alpha}\in\Gamma^{4}_{Y}(\mu,\nu)$ and
@@
In order to overcome the above drawbacks, we define a cost function for which the OT plan $\alpha\in\Gamma(\mu_{0},\mu_{1})$ only approximately fulfills $\pi^{1,3}_{\sharp}\alpha=\Delta_{\sharp}P_{Y}$:
@@
The same idea has been pursued in [2], where the authors rescaled the $y$-costs to obtain a blocktriangular map in the Knothe-Rosenblatt setting [37, 51] and similarly in [31]. Note that [31] was published on ArXiv after our first version [15] of the present paper.
@@
The following proposition shows that the issue raised in Example 9 is addressed by $W_{2,d_{\beta}}$.
@@
Assume that $\mu,\nu\in P_{2,Y}(\mathbb{R}^{d}\times\mathbb{R}^{m})$ and let $\mu_{n},\nu_{n}$ be empirical measures that converge weakly to $\mu,\nu$. Then for a sequence $\beta_{k}\to\infty$ there exists an increasing subsequence $n_{k}$ and optimal plans $\alpha_{n_{k}}\in\Gamma(\mu_{n_{k}},\nu_{n_{k}})$ for $W_{2,d_{\beta_{k}}}(\mu_{n_{k}},\nu_{n_{k}})$ such that $\alpha_{n_{k}}$ converges weakly to an optimal plan $\alpha\in\Gamma^{4}_{Y}(\mu,\nu)$ for $W_{2,Y}(\mu,\nu)$.
@@ 7 Bayesian flow matching
In this section, we combine the conditional Wasserstein distance with Bayesian flow matching. We start by briefly recalling flow matching and its OT variant. Then we turn to the conditional setting, where we describe a method from the literature, which we call "random" Bayesian flow matching and introduce our new OT Bayesian flow matching.
@@
The objective $L_{FM}$ can be also derived differently, with ideas inspired by the score matching framework [32, 57]. Then instead of regressing to the true velocity field at $x_{t}$, they show that regressing to it has the same gradients when one conditions at $x$ ("conditional" flow matching [39, Theorem 2]), which leads to the same loss formulation.
@@
Random Bayesian Flow Matching
@@
This yields the random Bayesian flow matching loss
$L_{Y,\text{FM}}(\theta)=\mathbb{E}_{(z,y,x)\sim P_{Z}\otimes P_{Y,X},t\sim U([0,1])}[\|v_{t}^{\theta}(y,x_{t})-(x-z)\|^{2}].$ (32)
Under the assumption $y_{i}\neq y_{j}$ for $i\neq j$ this matching coincides with the only admissible plan in the conditional Wasserstein distance. In general however, according to Example 9, this approach does not approximate OT plans as $X$ and $Z$ are essentially independent. Furthermore drawing a minibatch $\left((z_{i},y_{i},x_{i})\right)_{i=1}^{I}$ corresponds to a random coupling of the $z$ and $x$ for each class which explains the name random Bayesian flow matching.
@@
Given a minibatch $(\bm{z},\bm{y},\bm{x})=\left((z_{i},y_{i},x_{i})\right)_{i=1}^{I}$ sampled from the product distribution $P_{Z}\otimes P_{Y,X}$, we can calculate the optimal map $T$ and corresponding conditional coupling $\alpha$ between $(\bm{y},\bm{z})$ and $(\bm{y},\bm{x})$. The plan $\alpha$ by construction is only supported on $(y_{i},z_{i},y_{i},(\pi^{2}\circ T)(y_{i},z_{i})))$. Now let $(y,z,y,x)\sim\alpha$, then we have
@@
$L_{Y,\text{OT}}(\theta)=\mathbb{E}_{((y,z,y,x)\sim\alpha,t\sim U([0,1])}[\|v^{\theta}_{t}(y,x_{t})-(x-z)\|^{2}].$ (33)
@@ 8.1 Particle flow convergence
In this example, we minimize $W_{Y,d_{\beta}}(P_{Y,X},P_{Y,Z})$ for the empirical measures. We consider the particle flow, i.e., the flow from $(Y,Z)$ to $(Y,X)$ for empirical distributions which minimizes the objective $\mathcal{D}\left(\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i},x_{i}},\frac{1}{n}\sum_{i=1}^{n}\delta_{y_{i},z(t)_{i}}\right)$ for an appropriate distance $\mathcal{D}$, see e.g. [4]. More concretely we follow a particle flow path, i.e., a curve starting with $z(0)_{i}\sim\mathcal{N}(0,I)$ which fulfills
@@
for an appropriate scaling $\eta$ and given joint samples $(y_{i},x_{i})_{i=1}^{n}$. We choose $\mathcal{D}$ as an approximation of $W_{2,d_{\beta}}$ by rescaling $Y$ and using the Sinkorn divergence as the sample based distance measure [23, 21], where the "blur" parameter is chosen so small that it is close to the Wasserstein distance. This way we can numerically verify the convergence in Proposition 10. Note that there are no neural networks involved in this example.
We see in Fig. 1 that increasing $\beta$ indeed yields plans which transport no mass in $Y$-direction anymore, which has the consequence that the generated images fit the corresponding label. It can be seen that already for $\beta=5$ each row only has one type of digit.
@@ 8.2 GMM example
Here we use an experimental setup from [28]: Recall the Bayesian inverse problem setting in (1), i.e. $Y=f(X)+\Xi$ where choose $P_{X}$ to be a GMM in $\mathbb{R}^{5}$ with 10 mixture components, uniformly chosen means in $[-1,1]$ and standard deviation 0.1. We apply a linear diagonal forward model $f=(f_{i,j})_{i,j=1}^{5}\in\mathbb{R}^{5\times 5}$ with $f_{i,i}=\frac{0.1}{i+1}$ and zero components otherwise and choose $\Xi=\mathcal{N}(0,0.1)$ as a standard Gaussian distribution with standard deviation $0.1$. This yields a posterior measure $P_{X|Y=y}$ which is a also a GMM [28, Lemma 11]. Therefore we can sample and evaluate the true posterior as groundtruth. We train a random Bayesian flow matching model according to $L_{Y,\text{FM}}$(32) and our OT Bayesian flow matching according to $L_{Y,\text{OT}}$(33) with the python package POT [22] on a fixed dataset of size $10000$, where we choose the best model according to a validation set of size $2000$ until the validation loss converges. For both approaches we use the same feed-forward neural network which contains around 140k parameters. We evaluate them using the Sinkhorn distance with blur 0.05 [23] with the package GeomLoss [21] averaged with 100 posteriors and over 3 training and test runs with randomly sampled mixtures. The sampling is done via an explicit Euler discretization of 10 steps. Our proposed OT Bayesian flow matching model trained according to $L_{Y,\text{OT}}$ with $\beta=20$ obtains an average Sinkhorn distance of 0.0217, whereas the random Bayesian flow matching model obtains a value of 0.0278. In Figure 2 one can see that both models approximate the posterior very well.
Figure 2: Posterior histograms for different methods with random Bayesian flow matching on the left and our OT Bayesian flow matching on the right. Ground truth posterior is in orange and model prediction in blue.
@@ 8.3 Class conditional image generation
FID 4.97 4.94 4.38 4.33 4.10 4.92
Epochs 380 360 400 340 380 420
Figure 3: Class Conditional CIFAR results for different choices of $\beta$ and for the diagonal couplings. Additionally FID results are reported using an adaptive step size solver.
We apply our Bayesian OT flow matching for conditional image generation. We choose the condition $Y$ to be the class labels in order to generate samples of CIFAR10 for a given class. We simulate the flows for different values of $\beta$, by which we mean that we rescale the $Y$-component by $\beta$ as mentioned in Remark 11. We also simulate flows using the "diagonal" plans which coincide with the diagonal Bayesian flow matching objective [58]. For inference we simulate the flow ODE [17] using an adaptive step size solver (Runge-Kutta of order 5) . The samples in Fig. 3 are generated using the adaptive step size solver and sorted by class labels. For low values of $\beta$ we see that the resulting samples do not match their class labels, increasing $\beta$ leads to accurate class representations. The samples are generated given the labels of the training samples, therefore we see improved FID results as $\beta$ increases. The diagonal flow matching objective has the correct class representations, however since the associated couplings are not optimal our experiments suggest that this leads to higher variance during training and therefore slightly lower image quality, see [55] for more details on the advantages of OT based flow matching. We run each method for 500 epochs and compute an FID over the validation set every 20 epochs. Then for each method we choose the best checkpoint and report the results. The code is written in PyTorch [47] and is available online [^1] https://github.com/JChemseddine/Conditional_Wasserstein_Distances.